OpenStudy (anonymous):

3 + 4 log little x 4=5

OpenStudy (anonymous):

lol little x

OpenStudy (anonymous):

its underneath the log lol

OpenStudy (anonymous):

3 + 4log little x 4 = 5 4log little x 4 = 2 log little x 4 = 1/2 ...now it will become big x 4 = x^(1/2) hence x=16

OpenStudy (anonymous):

lol :D

OpenStudy (anonymous):

@ ishaan

OpenStudy (anonymous):

thanks